OpenStudy (anonymous):

solve using substitution 5x+3y=-34 -7x+y=58

OpenStudy (amistre64):

id rather use elimination

OpenStudy (anonymous):

yea elimination is simpler

OpenStudy (anonymous):

whatever works best for you

OpenStudy (amistre64):

but isnt that against the rules? it say we have to use subs

OpenStudy (anonymous):

either way you use, you will end up with the same answer........right?

OpenStudy (amistre64):

5x +3y =-34 -7(-3)x +y(-3) = 58(-3) --------------------- x(5+21) = 58(-3) -34 x = (58(-3) -34)/(5 + 21)

OpenStudy (amistre64):

yep, same answer as long as you do it right :)

OpenStudy (amistre64):

when you math it out; x = -8 i believe

OpenStudy (amistre64):

use that in either equation to determine the value of y

OpenStudy (anonymous):

does y=2

OpenStudy (amistre64):

ax + by = t ; (-c) cx + dy = z ; (a) -cax -cby = -tc cax +ady = za --------------- y(ad - cb) = za - tc y = (za-tc) / (ad - cb) ^^^ this part is known as a determinant of a matrix i think

OpenStudy (amistre64):

5(-8) +3y =-34 3y = -34 -5(-8) y = (-34 -5(-8))/3 = 2 yes

OpenStudy (anonymous):

Thanks !!!! :)

OpenStudy (amistre64):

youre welcome :)